### v2

# MRIQT: Physics-Aware Diffusion Model for Image Quality Transfer in Neonatal Ultra-Low-Field MRI

Abstract
Portable ultra-low-field MRI (uLF-MRI, T) offers accessible neuroimaging for neonatal care but suffers from low signal-to-noise ratio and poor diagnostic quality compared to high-field (HF) MRI. We propose MRIQT, a 3D conditional diffusion framework for image quality transfer (IQT) from uLF to HF MRI. MRIQT combines realistic K-space degradation for physics-consistent uLF simulation, -prediction with classifier-free guidance for stable image-to-image generation, and an SNR-weighted 3D perceptual loss for anatomical fidelity. The model denoises from a noised uLF input conditioned on the same scan, leveraging volumetric attention-UNet architecture for structure-preserving translation. Trained on a neonatal cohort with diverse pathologies, MRIQT surpasses recent GAN and CNN baselines in PSNR % with over the state of the art, while physicians rated of its outputs as good quality with clear pathology present. MRIQT enables high-fidelity, diffusion-based enhancement of portable ultra-low-field (uLF) MRI for reliable neonatal brain assessment. Implementation is available online111https://github.com/albarqounilab/MRIQT.
Index Terms— Portable Ultra-Low-Field MRI, Image Quality Transfer, 3D-Diffusion Models.
1 Introduction
Monitoring neonatal brain development is crucial for assessing early neurological function and detecting developmental disorders, particularly since the brain is highly vulnerable to injury [groteklaes4888105identifying]. Neonatal brain injuries and disorders can lead to death, over 2.3 million deaths within the first few days of life [who2024newborn], or long-term cognitive and motor impairments. Early detection of such pathologies is therefore essential to enable timely medical intervention, reduce infant mortality, and improve neuro-developmental outcomes. Neonatal neuroimaging aids in understanding that development.
In clinical practice, diagnosis typically begins with physical and neurological examinations, followed by cranial ultrasound (cUS) for infants exhibiting abnormal signs. While cUS is fast and bedside-accessible, it is highly operator-dependent and prone to artifacts, often yielding inconclusive results. High-field magnetic resonance imaging (HF-MRI) remains the gold standard for structural and functional assessment, providing high-resolution and high-contrast images of the developing brain. However, HF-MRI’s use in neonatal settings is severely constrained by the need for infant sedation, acoustic noise, prolonged scan times, limited portability, and high installation and maintenance costs, making it inaccessible in many neonatal intensive care units (NICUs).
Recently, ultra low-field MRI (uLF-MRI) has emerged as a portable, low-cost, and bedside-compatible alternative. It significantly reduces scan time and obviates the need for sedation. Yet, the dramatic reduction in magnetic field strength yields a much weaker signal-to-noise ratio, poor tissue contrast, and reduced spatial resolution, thereby limiting its diagnostic reliability. Bridging the quality gap between uLF and HF MRI, known as Image Quality Transfer (IQT), is thus crucial to unlocking the clinical potential of portable MRI systems for neonatal care. Enhancing such scans could facilitate the assessment process and aid in guiding intensive care therapy [sabir2023feasibility].
Deep learning-based super-resolution (SR) and image-to-image translation methods have shown promise for IQT by learning mappings from low- to high-quality MRI domains. To date, most uLF-to-HF IQT efforts have been GAN-based, such as LoHiResGAN [Islam2023improving], which employs a slice-wise PatchGAN discriminator [demir2018patch], and SFNet [Tap_SuperField_MICCAI2024], which integrates a Swin Transformer [liu2021swin] with dual-channel diffusion for uLF–HF data augmentation. Other approaches, including LF-SynthSR from the Freesurfer suite [iglesias2021joint, iglesias2023synthsr, iglesias2022quantitative], use convolutional networks for generating isotropic T1-weighted scans, while GAMBAS [baljer2025gambas] combines ConvNets with structured state-space blocks [gu2021efficiently] and Hilbert-curve serialization to model long-range spatial dependencies. Despite their effectiveness, GAN-based models often suffer from training instability, mode collapse, and poor structural fidelity, which are especially critical for clinical adoption. Moreover, most existing IQT models are trained on healthy adult or pediatric data, limiting their generalizability to neonatal pathologies.
Diffusion models, now the state-of-the-art in generative modeling, have recently demonstrated superior stability and fidelity in medical image synthesis. Prior works such as Med-DDPM [mobaidoctor2023medddpm], MU-Diff [dayarathna2025mu], and DiffusionIQT T [kim20233d] explored MRI generation and super-resolution, yet these methods typically rely on simulated low-field inputs generated via simple downsampling and blurring of HF images, which fail to capture the true physics of low-field signal formation. Furthermore, none have specifically addressed neonatal populations with real pathological variability.
In this work, we introduce MRIQT, a fully 3D conditional diffusion framework for ultra-low- to high-field brain MRI quality transfer. Our main contributions are as follows:
1.
We propose a fully 3D conditional diffusion model for volumetric IQT between uLF and HF MRI, designed to preserve anatomical structures while enhancing resolution and contrast.
2.
We train MRIQT on a neonatal-specific dataset encompassing diverse pathologies (e.g., hemorrhage, tumors, and structural malformations), enabling IQT for real-world NICU deployment.
3.
We derive a physics-aware K-space transfer function from paired uLF–HF scans to generate realistic synthetic uLF data from HF images, which enables unpaired training yet generalization to real uLF acquisitions without paired supervision.
4.
We introduce a 3D VGG-like perceptual loss that leverages feature-space similarity to guide diffusion sampling and early-stopping criteria, improving fine structural detail and convergence stability.
Together, these innovations make MRIQT the first diffusion-based framework tailored for clinically meaningful neonatal MRI quality transfer, bridging the fidelity gap between portable uLF and diagnostic HF scans, and moving toward practical, accessible neuroimaging for early-life care.
2 Methodology
We propose MRIQT, a 3D conditional diffusion framework for ultra-low-field to high-field (uLF–HF) MRI quality transfer. MRIQT combines (i) a physics-based K-space degradation model for realistic uLF simulation, (ii) a 3D conditional DDPM with classifier-free guidance (CFG) and -parametrization [salimans2022progressive], and (iii) a 3D perceptual feature loss for structure-preserving generation and adaptive early stopping.
2.1 Problem Setup
Given uLF and HF volumes , the goal is to learn a mapping that preserves anatomy from uLF while recovering HF-level resolution and contrast. Training uses both real and simulated uLF–HF pairs, where the latter are generated from HF data using a learned K-space transfer function.
2.2 K-space Simulation
For paired samples, we compute Fourier transforms and and estimate a complex transfer function by solving a Tikhonov-regularized least squares:
Synthetic uLF data are generated as . Matching radial power spectra confirms that simulated and real uLF scans share similar frequency behavior, enabling unpaired training on large HF datasets.
2.3 3D Conditional Diffusion
MRIQT extends the DDPM [ho2020denoising] to 3D volumes. The forward process adds Gaussian noise (Figure 2(a)), while the reverse model learns to predict and remove the noise conditioned on the uLF input. Instead of unconditional sampling, we use image-to-image diffusion, initializing denoising from a partially noised uLF scan, which preserves structural fidelity while allowing generative refinement (Figure 2(b)), as in [saharia2022image].
2.4 Classifier-Free Guidance
To balance uLF faithfulness and HF detail, we apply CFG [ho2022classifier]. During training, conditional and unconditional modes are alternated. At inference, predictions are merged as , where adjusts the strength of guidance and .
2.5 -Parametrization
Instead of predicting noise , we use the stable -prediction [salimans2022progressive], where . The model predicts , from which we recover . This yields smoother gradients across timesteps and sharper structural details than -prediction.
2.6 3D Perceptual Feature Extractor
Conventional 2D perceptual networks in MONAI [zhang2018unreasonable] fail on volumetric data. We therefore train a 3D VGG-like feature extractor on our neonatal HF dataset. The perceptual loss both (i) enhances contrast and fine texture during training and (ii) identifies an optimal sampling start step where uLF and HF become perceptually similar, improving efficiency and convergence.
2.7 Training Objective
The total loss combines -prediction with SNR-weighted perceptual alignment:
where denotes multi-scale feature maps from the 3D extractor, are layer weights, down-weights perceptual loss at high-noise steps, and is a regularization parameter.
3 Experiments and Results
| Method | PSNR | Pearson | SSIM | MS-SSIM | LPIPS | MAE |
| Baseline [uLF-HF] | ||||||
| LoHiResGAN [Islam2023improving] | ||||||
| LF-SynthSR [iglesias2021joint, iglesias2022quantitative] | ||||||
| SFNet [Tap_SuperField_MICCAI2024] | ||||||
| GAMBAS [baljer2025gambas] | ||||||
| MRIQT | ||||||
| -pred, | ||||||
| -pred, | ||||||
| -pred, (Ours) |
We evaluate MRIQT across quantitative, qualitative, and ablation studies to assess its ability to enhance uLF brain MRI quality and fidelity toward HF references. Experiments are designed to (i) validate the accuracy of our physics-based uLF simulation, (ii) benchmark MRIQT against state-of-the-art IQT and SR models, and (iii) isolate the contribution of key design components, including -prediction, perceptual loss, and CFG weighting.
3.1 Experimental Setup
3.1.1 Dataset
All data were collected at the Department of Neonatology and Pediatric Intensive Care Medicine, University Hospital Bonn, using both HF and uLF MRI scanners. T1-weighted uLF data were acquired using a 64 mT SWOOP (Hyperfine) scanner (1.51.52 mm anisotropic resolution). HF data were obtained from multiple clinical 3 T scanners with heterogeneous resolutions ranging from (0.350.340.43 mm) to (0.89 0.89 2 mm). The cohort included neonates under six months of age with nine different pathologies. After quality control, the dataset comprised 50 matched uLF–HF pairs, 100 HF-only scans, and 130 uLF-only scans. Matched pairs were used for K-space transfer estimation, determination, and final testing, while unpaired HF data were used for training and unpaired uLF data for reader study evaluation. All scans were converted from DICOM to NIfTI and organized following the BIDS standard [gorgolewski2016brain]. Preprocessing used ANTs [avants2011reproducible]: bias-field correction, resampling to 1 mm3, affine registration to the MNI152NLin2009cAsym template for 0–2 months [fonov2009unbiased], and cropping to a voxel grid.
3.1.2 Implementation Details
All experiments were conducted on an NVIDIA A100 (80 GB) GPU using PyTorch, with our implementation built upon the Med-DDPM framework [mobaidoctor2023medddpm]. Our 3D residual U-Net with attention bottleneck was trained for 30,000 epochs using AdamW with a weight decay, a 1,500-epoch warmup, and a cosine annealing scheduler (peak LR: ). Diffusion steps were ; inference using the learned step achieved a 35% speedup over standard sampling. The best CFG setup used and . The perceptual weight was fixed at . The 3D VGG-like feature extractor was trained via 5-fold cross-validation () using Adam and a linearly decaying LR schedule.
3.1.3 Baselines and Evaluation Metrics
We compare MRIQT against four state-of-the-art (SoTA) IQT and SR methods: LoHiResGAN [Islam2023improving], LF-SynthSR [iglesias2021joint, iglesias2022quantitative], SFNet [Tap_SuperField_MICCAI2024], and GAMBAS [baljer2025gambas]. Tests were conducted on 46 matched uLF–HF pairs (34 for LF-SynthSR, which requires T2-weighted input). Note that SFNet and GAMBAS were pre-trained on T2-weighted data, whereas MRIQT focuses on T1-weighted MRI.
We employed the following evaluation metrics; Pixel fidelity metrics: PSNR, MAE, RMSE, and Pearson correlation; and Perceptual and structural metrics: SSIM, MS-SSIM, and LPIPS. Both PSNR and correlation highlight reconstruction accuracy, while SSIM/MS-SSIM and LPIPS measure perceptual realism and structural fidelity. Statistical significance, w.r.t Ours, was assessed using a pair-wise t-test across the paired test sample at .
3.2 Quantitative Comparison
Table 1 summarizes the results on the paired subset (34 T1w samples with corresponding T2w scans). Conventional fidelity metrics such as PSNR, MAE, and SSIM primarily capture voxel-wise and structural intensity agreement rather than perceptual realism. Higher values indicate closer pixel correspondence; however, they do not necessarily correlate with anatomical fidelity or clinical interpretability as observed in the baseline results (first row of Table 1), where the high PSNR and SSIM do not correspond to superior visual quality. MRIQT achieves the highest PSNR (,dB) and competitive SSIM/MS-SSIM values compared to existing GAN- and CNN-based methods, reflecting improved noise suppression and detail restoration. Although SFNet attains slightly higher SSIM, its perceptual weighting tends to exaggerate details. In contrast, MRIQT achieves a lower MAE, supporting that diffusion-based denoising preserves anatomical texture while mitigating over-smoothing. The top-performing competitors in SSIM and MAE are not statistically superior to MRIQT. The Pearson correlation improvement over LF-SynthSR (+0.19) further highlights MRIQT’s robustness to intensity variations inherent to the uLF contrast domain. While GAMBAS achieves the highest correlation due to its global coherence focused objective, with L1-dominant loss (), this may compromise local structural fidelity that MRIQT explicitly preserves. Overall, MRIQT provides a balanced trade-off across all metrics, recovering perceptual image quality while maintaining both anatomical and global fidelity.
3.3 Ablation Studies
Replacing -prediction with -prediction decreases PSNR and LPIPS, confirming -prediction’s stability and fine-detail advantage. Removing the perceptual loss () leads to minor degradation in structure metrics (SSIM, MS-SSIM), showing that perceptual supervision primarily enhances contrast and local sharpness. The CFG hyperparameter sweep indicated that moderate guidance () balances realism and fidelity, avoiding hallucinated features. Overall, each contribution, -prediction, perceptual loss, and CFG, synergistically improves quantitative and perceptual outcomes.
3.4 Qualitative Analysis
Figure 3 compares uLF, SoTA, and MRIQT reconstructions. Competing GAN-based models (e.g., LoHiResGAN, SFNet) tend to oversharpen edges or hallucinate fine structures, while MRIQT consistently restores cortical boundaries and deep-gray matter contrast without introducing artifacts. MRIQT also preserves pathological features such as hemorrhagic lesions and ventricular enlargement, aligning closely with HF image. The denoising smoothness and texture recovery further confirm that diffusion-based IQT generalizes effectively across unseen pathologies.
3.5 Reader Study
Three raters, experienced in neonatal neuroimaging, independently evaluated generated scans by our model to assess the perceived image quality with grades; bad, acceptable, and good quality as , respectively. They also checked for the presence of any pathologies as a Quality Indicator. For pathology detection, we achieved an average accuracy of and a sensitivity of against the ground truth labels, with an average image quality rating of across raters. The raters were not exposed to the original scans; hence, this study is not clinically applicable, as the raters were not provided with the original scans to assess the diagnostic impact. More extensive paired studies are planned for the future.
4 Conclusion
We presented MRIQT, a 3D conditional diffusion framework for image quality transfer between uLF and HF brain MRI. Designed to meet the clinical need for reliable and portable neonatal imaging, MRIQT combines a physics-aware K-space simulation, -prediction, classifier-free guidance, and a 3D perceptual loss to achieve structure-preserving enhancement of uLF scans.
Experiments on neonatal data show that MRIQT surpasses existing GAN- and CNN-based baselines in PSNR and perceptual quality, maintaining high anatomical fidelity without hallucinations. Reader studies further confirm improved interpretability and diagnostic plausibility of MRIQT reconstructions.
Future work will extend MRIQT to multimodal and real-time reconstruction, enabling high-quality, bedside neuroimaging in neonatal intensive care and resource-limited settings.
Compliance with ethical standards. This study was performed in line with the principles of the Declaration of Helsinki. Approval was granted by the Ethics Committee of UniBonn (Ethics Nr. 167/22).